# An Efficient Scene Coordinate Encoding and Relocalization Method

### v1

@@ Abstract
Scene Coordinate Regression (SCR) is a visual localization technique that utilizes deep neural networks (DNN) to directly regress 2D-3D correspondences for camera pose estimation. However, current SCR methods often face challenges in handling repetitive textures and meaningless areas due to their reliance on implicit triangulation. In this paper, we propose an efficient scene coordinate encoding and relocalization method. Compared with the existing SCR methods, we design a unified architecture for both scene encoding and salient keypoint detection, enabling our system to focus on encoding informative regions, thereby significantly enhancing efficiency. Additionally, we introduce a mechanism that leverages sequential information during both map encoding and relocalization, which strengthens implicit triangulation, particularly in repetitive texture environments. Comprehensive experiments conducted across indoor and outdoor datasets demonstrate that the proposed system outperforms other state-of-the-art (SOTA) SCR methods. Our single-frame relocalization mode improves the recall rate of our baseline by 6.4% and increases the running speed from 56 to 90. Furthermore, our sequence-based mode increases the recall rate by 11% while maintaining the original efficiency.
@@ I Introduction
Visual relocalization is crucial for many applications such as mobile robots and augmented reality, offering a practical balance between cost and accuracy [1]. Its core task involves estimating the camera pose using an existing map, which can be broadly divided into two categories: explicit maps and implicit maps. Explicit maps consist of a 3D point cloud combined with visual descriptors of 2D features [2, 3], while implicit maps encode the scene using a neural network [1, 4]. Feature matching (FM)-based methods construct explicit maps using techniques like structure-from-motion (SFM) [3, 5] or simultaneous localization and mapping (SLAM) [2, 6]. During relocalization, these methods rely on image retrieval and feature matching to establish 2D-3D correspondences between pixels in the query image and 3D points in the map. However, this approach often suffers from large map sizes due to the storage of visual features and 3D models. Additionally, the process can be computationally inefficient, as the feature matching needs to be performed between the query image and multiple candidates.
In contrast, scene coordinate regression (SCR) methods encode map information into a neural network [1, 7]. They directly regress the 3D coordinates in the world coordinate system correspondences to the image pixels, eliminating the need for explicit 2D-3D corresponding searching. Compared to FM-based methods, SCR methods enable faster relocalization with a smaller map. Recently, ACE [8] proposes a generalized convolutional neural network (CNN) to extract feature maps for scene encoding. Building on this pre-trained backbone, it utilizes a compact 4MB-sized multilayer perceptron (MLP) head as the map, enabling new scene encoding in five minutes only without requiring a 3D model or depth information. This approach achieves state-of-the-art accuracy and efficiency.
@@
Despite their advantages, SCR methods face several notable limitations. First, the presence of many texture-less and non-informative regions in the image significantly increases the map encoding errors [9]. Second, existing RGB-only SCR methods rely heavily on implicit triangulation, i.e., the neural network must consistently identify and map the same regions across different images to accurate 3D points. This reliance causes a sharp decline in performance in environments with repetitive textures [10]. Third, during relocalization, all pixel-to-3D matches are considered for pose estimation since the reliability of specific regions is unknown, which significantly reduces efficiency. Efforts have been made to address these challenges. FocusTune [9] mitigates encoding errors by focusing on informative regions but requires a 3D model of the scene, which limits its practicality. For the second issue, GLACE [10] introduces global image descriptors to better distinguish similar regions across images, yet this approach comes at the cost of reduced mapping and localization efficiency.
@@ II-A Feature Matching Based Relocalization
However, FM-based methods require storing 2D keypoints, visual descriptors, and a 3D point cloud, resulting in relatively large map sizes. For large-scale scenes, managing these maps demands significant memory resources. Furthermore, during relocalization, multiple candidate frames are typically retrieved for feature matching, which imposes computational overhead and reduces efficiency.
@@ II-B Absolute Pose Regression
Absolute pose regression (APR) approaches use a DNN to regress camera poses from RGB images directly. They have also garnered notable attention recently due to their good efficiency and simple architecture. PoseNet [15] is the first ARP solution which uses GoogLeNet [16] as the backbone. The following works explore different architectural designs to enhance ARP systems [17, 18]. Other works leverage sequential information to improve APR accuracy by incorporating temporal filtering [19] or visual odometry [20]. Recent ARP works utilize novel synthesis techniques to generate high-quality training data [21] or integrate them for fine-tuning [22]. However, ARP approaches still fall short in accuracy compared to SCR and FM-based methods.
@@ II-C Scene Coordinate Regression
Different from ARP approaches, SCR methods regress the corresponding 3D coordinates in the scene for 2D pixels in the query image. Most existing SCR methods require depth information or a 3D model of the scene. Dong et al. [7] utilize a pre-learned feature extractor and scene region classifier to encode a scene with few-shot images, achieving efficient mapping. However, the additional modules, e.g., feature detection and hierarchical classification, limit its localization speed to only $5$ even on a powerful GPU. Zhou et al. [23] propose KFNet. They leverage the sequential information by integrating the traditional Kalman filter [24] into the SCR pipeline. Bui et al. [25] propose D2S, which builds on a pre-trained sparse feature extraction network. They establish a mapping between feature descriptors and 3D scene points.
DSAC++ [26] is the first SCR method that does not rely on depth or 3D model information. However, it requires a long training time. DSAC* [1] simplifies the DSAC++ training procedure by unifying two separate initialization steps, reducing the training time from 6 days to 15 hours. ACE [8] replaces the backbone of DSAC* with a pre-trained CNN, requiring only a new header to be trained when encoding a new scene, which greatly improves the mapping efficiency. This reduces the mapping time to 4 minutes and the map size to 4MB. FocusTune [9] and GLACE [10] are variations of ACE. FocusTune projects a 3D model onto training images to avoid encoding meaningless image regions, but this introduces the dependency on the 3D model. GLACE incorporates global image descriptors into the ACE pipeline to enhance the distinguishability of similar regions.
@@ III-A Problem Statement
The objective is to estimate the pose $\mathbf{T}_{wc}$ of a given image $I$, where $\mathbf{T}_{wc}$ denotes the transformation from the camera coordinate system to the world coordinate system. Given a set of 2D-3D correspondences $\mathcal{C}={(\mathbf{p}_{i},\mathbf{P}_{i}^{w})}$, where $\mathbf{p}_{i}$ represents a pixel in the image $I$ and $\mathbf{P}_{i}^{w}$ is its corresponding 3D point in the world coordinate system, the pose can be efficiently computed using PnP algorithms or iterative optimization techniques. To obtain these correspondences $\mathcal{C}$, SCR methods aim to learn a mapping from 2D pixels to their corresponding 3D points:
@@ IV-A Overview
The overall structure of the proposed system is depicted in Fig. 2. It comprises two main components: a scene-agnostic module and a scene-specific module. The scene-agnostic module generates feature vectors and saliency scores for image patches, while the scene-specific module, designed as an MLP head with several 1D convolutional layers, maps these feature vectors to 3D scene points. The detailed design of the scene-agnostic module is presented in Section IV-B. Section IV-C describes how sequential information is utilized to enhance the training of the scene-specific module, and Section IV-D outlines our relocalization methods.
@@ IV-B Scene-Agnostic Module
In this section, we aim to design an efficient and scene-agnostic module capable of (1) detecting informative image patches and (2) mapping these patches into high-dimensional features for scene encoding. As many learning-based patch and keypoint detection methods have been proposed [3, 27], a straightforward solution is to utilize them directly for salient region detection along with the ACE backbone. However, employing two separate deep neural networks would significantly reduce online efficiency. Inspired by [3, 6], we think that these two tasks can be accomplished using a shared backbone. To achieve this, we extend the ACE backbone by adding a lightweight keypoint detection head (KDH). The KDH consists of just three convolutional layers, introducing trivial computational overhead. Its input is the feature map from the backbone, and its output is a tensor $\mathcal{T}_{K}\in\mathbb{R}^{65\times\frac{H}{8}\times\frac{W}{8}}$, where $H$ and $W$ denote the height and width of the input image, respectively. The 65 channels correspond to an $8\times 8$ grid region and a dustbin channel representing the absence of keypoints. This tensor is processed through a softmax layer and resized into a $H\times W$ heatmap, which scores the saliency of the corresponding image pixels. Since the KDH, like the backbone $f_{B}$ in Eq. 2, is also scene-agnostic, we combine them into a unified module denoted as $f_{S}$, with its parameters represented by $\mathbf{w}_{S}$. Then the second equation in Eq. 2 is updated to:
@@
We train $f_{S}$ using a two-step pipeline on the ScanNet dataset [28]. In the first step, we follow the approach outlined in [8] to train the backbone $f_{B}$. In the second step, the parameters of $f_{B}$ are fixed, and we utilize distillation techniques to train the KDH. Theoretically, any keypoint detector could serve as the teacher network. For our setup, we use the prediction of SuperPoint [27] as supervision signals. The loss function is defined as:
@@
where $\mathcal{L}_{CE}$ is the cross-entropy loss and $\mathcal{T}^{\prime}_{K}$ is the prediction of the teacher network. We also experimented with a multi-task learning approach [3] to train $f_{S}$, but found that the two-step pipeline produced better results. This is probably because the difficulty of the two tasks differs significantly: specifically, training a keypoint detector is much simpler than training a universal scene encoder. This disparity makes it challenging to balance the two tasks effectively.
Figure 3: Sampling strategies comparison. ACE randomly samples image patches. FocusTune performs uniform sampling around the projected points of the 3D model. We use keypoint detection to select repeatable image patches.
@@ IV-C Sequence-Based Mapping
Buffer sampling is the key to fast mapping. Since the backbone, i.e., $f_{B}$ or $f_{S}$, is scene-agnostic, it only needs to be executed once when mapping a new scene. The resulting outputs, $\{\mathbf{f}_{i}\}$, can be cached and subsequently used to train $f_{H}$. To enhance efficiency, not all image patches are utilized during training. In ACE [8], 1024 image patches are randomly sampled per image, and their corresponding $\mathbf{f}_{i}$, $\mathbf{p}_{i}$, camera intrinsics, and camera poses are stored in a training buffer. However, this strategy may sample many meaningless regions. To address this, FocusTune [9] projects the 3D point cloud of the scene onto the training images and performs uniform sampling around the projected points. While this approach better targets informative areas, it cannot guarantee consistent sampling across multiple images.
In contrast, we utilize the keypoint detection results from Section IV-B to select the top 1,000 patches with the highest confidence scores in each image, aligning with the number used by ACE. This approach ensures consistent patch selection across image sequences. It also ensures the same patches can be used during mapping and relocalization. We show a comparison of patch selection in Fig. 3.
@@
For each image patch, $\mathcal{L}_{S,i}$ uses information from a single image to supervise the prediction of $\mathbf{P}_{i}^{w}$ for each $\mathbf{f}_{i}$, so we call it self-loss. However, in classical geometry, a single viewpoint constrains a scene point to lie along a line of sight rather than determining its precise 3D position. For instance, $\mathcal{L}_{S,i}$ would be zero if the predicted $\mathbf{P}_{i}^{w}$ falls anywhere on the inverse-projection line of $\mathbf{p}_{i}$. Therefore, SCR systems heavily rely on effective implicit triangulation, i.e., the neural network must consistently identify the same region across multiple images so that the scene point can be constrained by observations from different viewpoints. Achieving this is particularly challenging in complex or repetitive texture environments, as shown in [10].
@@
For each image patch, we store the corresponding feature vector $\mathbf{f}_{i}$, the pixel $\mathbf{p}_{i}$, the camera intrinsic $\mathbf{K}$, the camera pose $\mathbf{T}_{wc}$, the pixel matched on the keyframe $\mathbf{p}_{k,j}$, and the keyframe pose $\mathbf{T}_{wk}$, in the training buffer. We shuffle these patches and select $\mathcal{N}_{b}=5120$ patches in each iteration to compute the loss:
@@
IV-D Dual-Mode Relocalization
In this section, we introduce how to reuse the map for relocalization. Our system supports two relocalization modes. The first is single-mode, which estimates the camera pose using a single image. This mode is efficient and straightforward. As shown in Fig. 2, the framework generates a set of 2D keypoints and their corresponding predicted 3D scene points. Using these 2D-3D correspondences, we apply PnP and RANSAC algorithms to compute the optimal pose by maximizing the number of inliers. Like most existing SCR methods, our single-mode does not leverage the constraints between adjacent frames in an image sequence. However, in real-world robotic applications, the previous frame provides valuable prior information for pose estimation. Therefore, we also design a sequence-mode that utilizes sequential information to enhance the relocalization. The details of this mode are introduced in the following section.
@@ IV-D Dual-Mode Relocalization
where $\mathfrak{g}$ represent the PnP solver with RANSAC.
@@
The pose estimation in Eq. 10 only uses 2D-3D correspondences of historical frames and feature tracking results, so $\mathbf{\bar{T}}_{wn}$ can be refined with new predictions from the network for $\mathcal{I}_{n}$. Let $\mathcal{\hat{X}}_{n}$ and $\mathcal{\hat{Z}}_{n}$ denote the newly detected keypoints and the corresponding 3D scene points predicted by the network. Since the set of 2D-3D correspondences $\{\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\}$ typically contains many outliers, we first use PnP with RANSAC to compute a pose $\mathbf{\hat{T}}_{wn}=\mathfrak{g}\left(\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\right)$, then fuse it with $\mathbf{\bar{T}}_{wn}$ at the pose level, rather than directly using a Kalman filter. We perform a simple weighted average on the manifold to obtain the updated pose:
@@
The maintained scene point set also requires updating based on new observations from $\mathcal{I}_{n}$. Besides adding newly detected points and their corresponding predicted 3D points to $\{\mathcal{\bar{X}}_{n},\mathcal{\bar{Y}}_{n}\}$, we also update tracked scene points. For a scene point $\mathbf{\bar{P}}^{w}_{i}\in\mathcal{\bar{Y}}_{n}$ and its corresponding tracked 2D point $\mathbf{\bar{p}}_{i}^{n}\in\mathcal{\bar{X}}_{n}$, if the pair $(\mathbf{\bar{p}}_{i}^{n},\mathbf{\bar{P}}^{w}_{i})$ is an outlier in Eq. 10, it will be removed from $\mathcal{\bar{Y}}_{n}$ and $\mathcal{\bar{X}}_{n}$. Otherwise, $\mathbf{\bar{P}}^{w}_{i}$ is updated in a direction that reduces the reprojection error on $\mathcal{I}_{n}$. The update direction is illustrated in Fig. 4, where $\scriptstyle{P}$ and $\scriptstyle{p}$ represent $\mathbf{\bar{P}}^{w}_{i}$ and $\mathbf{\bar{p}}_{i}^{n}$ for simplicity. The inverse projection line of $\mathbf{\bar{p}}_{i}^{n}$ is denoted as $\scriptstyle{\overrightarrow{ND}}$, with $\scriptstyle{D}$ as the foot of the perpendicular. The coordinates of $\scriptstyle{D}$ are:

### v2

# Enhancing Scene Coordinate Regression with Efficient Keypoint Detection and Sequential Information

@@ Abstract
Scene Coordinate Regression (SCR) is a visual localization technique that utilizes deep neural networks (DNN) to directly regress 2D-3D correspondences for camera pose estimation. However, current SCR methods often face challenges in handling repetitive textures and meaningless areas due to their reliance on implicit triangulation. In this paper, we propose an efficient and accurate SCR system. Compared to existing SCR methods, we propose a unified architecture for both scene encoding and salient keypoint detection, allowing our system to prioritize the encoding of informative regions. This design significantly improves computational efficiency. Additionally, we introduce a mechanism that utilizes sequential information during both mapping and relocalization. The proposed method enhances the implicit triangulation, especially in environments with repetitive textures. Comprehensive experiments conducted across indoor and outdoor datasets demonstrate that the proposed system outperforms state-of-the-art (SOTA) SCR methods. Our single-frame relocalization mode improves the recall rate of our baseline by 6.4% and increases the running speed from 56 to 90. Furthermore, our sequence-based mode increases the recall rate by 11% while maintaining the original efficiency.
@@ I Introduction
Visual relocalization is crucial for many applications such as mobile robots and augmented reality, offering a practical balance between cost and accuracy [1]. The core task involves estimating the camera pose using a pre-built map. The map can be generally classified into two categories: explicit maps and implicit maps. Explicit maps consist of a 3D point cloud combined with visual descriptors of 2D features [2, 3], while implicit maps encode the scene using a neural network [1, 4]. Feature matching (FM)-based methods construct explicit maps using techniques like structure-from-motion (SFM) [3, 5] or simultaneous localization and mapping (SLAM) [2, 6]. During relocalization, these methods rely on image retrieval and feature matching to establish 2D-3D correspondences between pixels in the query image and 3D points in the map. However, this approach often suffers from large map sizes due to the storage of visual features and 3D models. Additionally, the process can be computationally inefficient, as the feature matching needs to be performed between the query image and multiple candidates.
In contrast, scene coordinate regression (SCR) methods encode map information into a neural network [1, 7]. They directly regress the 3D coordinates in the world coordinate system correspondences to the image pixels, eliminating the need for explicit 2D-3D corresponding searching. Compared to FM-based methods, SCR methods enable faster relocalization with a smaller map. Recently, ACE [8] proposes a generalized convolutional neural network (CNN) to extract feature maps for scene encoding. Building on this pre-trained backbone, ACE utilizes a compact 4MB-sized multilayer perceptron (MLP) head as the map, enabling new scene encoding in five minutes only without requiring a 3D model or depth information. This approach achieves state-of-the-art accuracy and efficiency.
@@
Despite their advantages, SCR methods still face several notable limitations. First, the presence of many texture-less and non-informative regions in the image significantly increases the map encoding errors [9]. Second, existing RGB-only SCR methods rely heavily on implicit triangulation, i.e., the neural network must consistently identify and map the same regions across different images to accurate 3D points. This reliance causes a sharp decline in performance in environments with repetitive textures [10]. Third, during relocalization, all pixel-to-3D matches are considered for pose estimation since the reliability of specific regions is unknown, which significantly reduces efficiency. Efforts have been made to address these challenges. FocusTune [9] mitigates encoding errors by focusing on informative regions but requires a 3D model of the scene, which limits its practicality. For the second issue, GLACE [10] introduces global image descriptors to better distinguish similar regions across images, yet this approach comes at the cost of reduced mapping and localization efficiency.
@@ II-B Absolute Pose Regression
Unlike FM-based methods, absolute pose regression (APR) approaches use a DNN to regress camera poses from RGB images directly. They have also garnered notable attention recently due to their good efficiency and simple architecture. PoseNet [15] is the first APR system. The subsequent works explore different architectural designs [16] or leverage sequential information [17] to enhance APR systems. Recent APR works utilize novel synthesis techniques to generate high-quality training data [18] or integrate them for fine-tuning [19]. However, APR approaches still fall short in accuracy compared to SCR and FM-based methods.
@@ II-C Scene Coordinate Regression
Different from FM-based methods and APR approaches, SCR methods regress the corresponding 3D coordinates in the scene for 2D pixels in the query image. Most existing SCR methods require depth information or a 3D model of the scene [20, 21]. Dong et al. [7] use a pre-trained feature extractor and scene classifier for efficient few-shot mapping, but extra modules limit localization speed to $5$ even on a high-end GPU. Zhou et al. [22] introduce KFNet, integrating a Kalman filter [23] into the SCR pipeline to leverage sequential information. Bui et al. [24] propose D2S, which maps sparse feature descriptors to 3D scene points using a pre-trained network.
DSAC++ [25] is the first SCR method independent of depth or 3D models but requires long training. DSAC* [1] streamlines DSAC++ by unifying initialization steps, cutting training time from 6 days to 15 hours. ACE [8] further improves efficiency by using a pre-trained CNN, reducing mapping time to 4 minutes and map size to 4MB. FocusTune [9], GLACE [10], and EGFS [26] are all variants of ACE, each introducing unique enhancements. FocusTune refines ACE by projecting a 3D model onto images to filter irrelevant regions, but it depends on the 3D model. EGFS refines region selection through an iterative process using a segmentation model and reprojection errors. Meanwhile, GLACE enhances ACE by incorporating global image descriptors for better region distinction. Our system also builds on ACE. However, unlike these methods, we utilize keypoint detection to filter out uninformative regions and leverage sequential information to reduce visual aliasing.
II-D Keypoint Detection
Keypoint detection is essential for visual localization. Traditional methods like FAST [27] and ORB [28] rely on handcrafted detectors and descriptors, offering efficiency but often lacking robustness. In contrast, learning-based approaches, such as SuperPoint [29] and R2D2 [30], leverage deep networks to enhance repeatability and precision. Recently, detector-free methods like D2-Net [31] jointly learn detection and description, improving performance in challenging scenarios. To optimize efficiency, keypoint detection can leverage a shared backbone network in multi-task learning, enabling integration with tasks like local-global feature extraction [3] or joint point-line detection [6]. Inspired by [3, 6], we use a shared backbone for both SCR and keypoint detection, improving system efficiency.
@@ III-A Problem Statement
The objective is to estimate the pose $\mathbf{T}_{wc}$ of a given image $I$, where $\mathbf{T}_{wc}$ denotes the transformation from the camera coordinate system to the world coordinate system. Given a set of 2D-3D correspondences $\mathcal{C}={(\mathbf{p}_{i},\mathbf{P}_{i}^{w})}$, where $\mathbf{p}_{i}$ represents a pixel in the image $I$ and $\mathbf{P}_{i}^{w}$ is the corresponding 3D point in the world coordinate system, the pose can be efficiently computed using PnP algorithms or iterative optimization techniques. To obtain these correspondences $\mathcal{C}$, SCR methods aim to learn a mapping from 2D pixels to their corresponding 3D points:
@@ IV-A Overview
The overall structure of the proposed system is depicted in Fig. 2. Our system consists of a scene-agnostic module and a scene-specific module. The scene-agnostic module generates feature vectors and saliency scores for image patches, while the scene-specific module, designed as an MLP head with several 1D convolutional layers, maps these feature vectors to 3D scene points. The detailed design of the scene-agnostic module is presented in Section IV-B. Section IV-C describes how sequential information is utilized to enhance the training of the scene-specific module, and Section IV-D outlines our relocalization methods.
@@ IV-B Scene-Agnostic Module
In this section, we aim to design an efficient and scene-agnostic module capable of (1) detecting informative image patches and (2) mapping these patches into high-dimensional features for scene encoding. As many learning-based patch and keypoint detection methods have been proposed [3, 29], a straightforward solution is to utilize them directly for salient region detection along with the ACE backbone. However, employing two separate deep neural networks would significantly reduce online efficiency. Inspired by [3, 6], we accomplish these two tasks with a shared backbone. To achieve this, we extend the ACE backbone by adding a lightweight keypoint detection head (KDH). The KDH consists of just three convolutional layers, introducing trivial computational overhead. The KDH takes the feature map from the backbone as input and outputs a tensor $\mathcal{T}_{K}\in\mathbb{R}^{65\times\frac{H}{8}\times\frac{W}{8}}$, where $H$ and $W$ denote the height and width of the input image, respectively. The 65 channels correspond to an $8\times 8$ grid region and a dustbin channel representing the absence of keypoints. This tensor is processed through a softmax layer and resized into a $H\times W$ heatmap, which scores the saliency of the corresponding image pixels. Since the KDH, similar to the backbone $f_{B}$ in Eq. 2, is scene-agnostic, we integrate them into a unified module, denoted as $f_{S}$, with parameters represented by $\mathbf{w}_{S}$. Then the second equation in Eq. 2 is updated to:
@@
We train $f_{S}$ using a two-step pipeline on the ScanNet dataset [32]. In the first step, we follow the approach outlined in [8] to train the backbone $f_{B}$. In the second step, the parameters of $f_{B}$ are fixed, and the distillation techniques are utilized to train the KDH. We use the prediction of SuperPoint [29] as supervision signals as it is simple, efficient, and performs excellently. The loss function is defined as:
@@
where $\mathcal{L}_{CE}$ is the cross-entropy loss and $\mathcal{T}^{\prime}_{K}$ is the prediction of the teacher network. We also experimented with a multi-task learning approach [3] to train $f_{S}$, but found that the two-step pipeline produced better results. This may be attributed to the substantial difference in difficulty between these two tasks. Specifically, training a keypoint detector converges much faster than training a universal scene encoder. This disparity in task difficulty may lead to suboptimal training within a multitask network [33].
Figure 3: Sampling strategies comparison. ACE [8] randomly samples image patches, while FocusTune [9] performs uniform sampling around the projected points of the 3D model. In contrast, we employ keypoint detection to identify repeatable image patches.
@@ IV-C Sequence-Based Mapping
Buffer sampling is the key to fast mapping. Since the backbone, i.e., $f_{B}$ or $f_{S}$, is scene-agnostic, it only needs to be executed once when mapping a new scene. The resulting outputs, $\{\mathbf{f}_{i}\}$, can be cached and subsequently used to train $f_{H}$. To enhance efficiency, not all image patches are utilized during training. In ACE [8], 1024 image patches are randomly sampled per image, and their corresponding $\mathbf{f}_{i}$, $\mathbf{p}_{i}$, camera intrinsics, and camera poses are stored in a training buffer. However, this strategy may sample many meaningless regions. To address this, FocusTune [9] projects the 3D point cloud of the scene onto the training images and performs uniform sampling around the projected points. While this approach better targets informative areas, FocusTune cannot guarantee consistent sampling across multiple images.
In contrast, we utilize the keypoint detection results from Section IV-B to select the top 1,000 patches with the highest confidence scores in each image, aligning with the number used by ACE. This approach guarantees consistent patch selection across image sequences, and also ensures the same patches are utilized during both mapping and relocalization. We show a comparison of patch selection in Fig. 3.
@@
For each image patch, $\mathcal{L}_{S,i}$ supervises the prediction of $\mathbf{P}_{i}^{w}$ for each $\mathbf{f}_{i}$ using only a single image, which is why we refer to it as self-loss. However, in classical multi-view geometry, triangulating a point successfully requires at least two observations [10, 34]. Therefore, SCR systems heavily rely on implicit triangulation, implying that the neural network must consistently recognize the same region across multiple images to ensure that scene points are constrained by observations from different viewpoints. This is particularly challenging under complex environments.
@@
For each image patch, we store the corresponding feature vector $\mathbf{f}_{i}$, the pixel $\mathbf{p}_{i}$, the camera intrinsic $\mathbf{K}$, the camera pose $\mathbf{T}_{wc}$, the pixel matched on the keyframe $\mathbf{p}_{k,j}$, and the keyframe pose $\mathbf{T}_{wk}$, in the training buffer. We follow ACE [8] to shuffle these patches and select $\mathcal{N}_{b}=5120$ patches in each iteration to compute the loss:
@@
IV-D Relocalization
IV-D1 Single-Mode
Our system supports two relocalization modes. The single-mode estimates the camera pose from a single image, offering efficiency and simplicity. As shown in Fig. 2, it detects 2D keypoints, predicts corresponding 3D scene points, and applies PnP with RANSAC to optimize the pose by maximizing inliers. Note that this mode also uses the map built with our sequence-based mapping method.
IV-D2 Sequence-Mode
Like most existing SCR methods, our single-mode does not leverage the constraints between adjacent frames in an image sequence during relocalization. However, in real-world robotic applications, the previous frame provides valuable prior information for pose estimation. Therefore, we also design a sequence-mode that utilizes sequential information to enhance the relocalization. Details of this mode are in the following section.
@@ IV-D2 Sequence-Mode
where $\mathfrak{g}$ represents the PnP solver with RANSAC.
@@
The pose estimation in Eq. 10 only uses 2D-3D correspondences of historical frames and feature tracking results, so $\mathbf{\bar{T}}_{wn}$ can be refined with new predictions from the network for $\mathcal{I}_{n}$. Let $\mathcal{\hat{X}}_{n}$ and $\mathcal{\hat{Z}}_{n}$ denote the newly detected keypoints and the corresponding 3D scene points predicted by the network. Since the set of 2D-3D correspondences $\{\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\}$ typically contains many outliers, we first use PnP with RANSAC to compute a pose $\mathbf{\hat{T}}_{wn}=\mathfrak{g}\left(\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\right)$, then fuse $\mathbf{\hat{T}}_{wn}$ with $\mathbf{\bar{T}}_{wn}$ at the pose level, rather than directly using a Kalman filter. We perform a simple weighted average on the manifold to obtain the updated pose:
@@
The maintained scene point set also requires updating based on new observations from $\mathcal{I}_{n}$. Besides adding newly detected points and their corresponding predicted 3D points to $\{\mathcal{\bar{X}}_{n},\mathcal{\bar{Y}}_{n}\}$, we also update tracked scene points. For a scene point $\mathbf{\bar{P}}^{w}_{i}\in\mathcal{\bar{Y}}_{n}$ and the corresponding tracked 2D point $\mathbf{\bar{p}}_{i}^{n}\in\mathcal{\bar{X}}_{n}$, if the pair $(\mathbf{\bar{p}}_{i}^{n},\mathbf{\bar{P}}^{w}_{i})$ is an outlier in Eq. 10, it will be removed from $\mathcal{\bar{Y}}_{n}$ and $\mathcal{\bar{X}}_{n}$. Otherwise, $\mathbf{\bar{P}}^{w}_{i}$ is updated in a direction that reduces the reprojection error on $\mathcal{I}_{n}$. The update direction is illustrated in Fig. 4, where $\scriptstyle{P}$ and $\scriptstyle{p}$ represent $\mathbf{\bar{P}}^{w}_{i}$ and $\mathbf{\bar{p}}_{i}^{n}$ for simplicity. The inverse projection line of $\mathbf{\bar{p}}_{i}^{n}$ is denoted as $\scriptstyle{\overrightarrow{ND}}$, with $\scriptstyle{D}$ as the foot of the perpendicular. The coordinates of $\scriptstyle{D}$ are: